OpenStudy (ttop0816):

am i correct?? x + y = 0 x - y = 8 What is the solution of the system shown? (-8, 12) (8, -12) <-- i chose this! (8, 12)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

8-(-12) is 8 +12 which would be 20, so x - y = 8 doesn't work (so no)

OpenStudy (ttop0816):

sorry! the question got cut off 1/2x+1/3y=0 1/4x-1/2y=8

OpenStudy (anonymous):

ya

OpenStudy (ttop0816):

@baby1234567 could you help me with this question too?? Going against the current, a boat takes 6 hours to make a 120-mile trip. When the boat travels with the current on the return trip, it takes 5 hours. What is the rate of the boat in still water? 2 mph 20 mph 22 mph <--- i think its this one... but im not sure ):

OpenStudy (anonymous):

120/6, so 20mph

OpenStudy (ttop0816):

i chose 20mph and it was incorrect...

OpenStudy (anonymous):

ya that's right

OpenStudy (anonymous):

oh crap, 22 was right, I thought it said it took 6 hours to make a 120 while in still. Sorry :/

OpenStudy (ttop0816):

@theloaad oh its fine! thank you still for trying to help me (: @baby1234567 sorry to keep asking but could you check if my answer for this question is correct?! A 20% acid solution is mixed with a 70% acid solution to get 50 liters of a 40% solution. Which of the following expressions is equal to 4(50)? 0.2x + 0.7y 2x + 7y <--- chose thisss x + y

OpenStudy (anonymous):

ya your right

OpenStudy (ttop0816):

@baby1234567 omg thank you a billionnnn you just saved me :D

OpenStudy (anonymous):

your welcome lol